# Osmolyte-Modulated Differential Capacitance and Disjoining Pressure for Nanoconfined Electrolytes: A Modified Poisson-Boltzmann Theory

### v1

@@ III.2 Disjoining pressure
Recent experimental studies by Govrin et al. [49] reveal that incorporating osmolytes into an electrolyte solution diminishes the electrostatic screening of charged surfaces by the mobile ions present in the solution, thereby amplifying the repulsive forces between them. In the context of slit-like pores with charged walls, this phenomenon is expected to result in an enhancement of the electrostatic component of the disjoining pressure as the concentration of zwitterionic molecules in the bulk solution increases. The MPB theory predicts this behavior, as illustrated in Figures 7 and 8. It suggests that the disjoining pressure will rise with increasing glycine concentration, applicable to both its peptide and dipeptide forms. Notably, this effect is observed at two distinct ranges of surface potentials: at low surface potentials, where the glycine concentration within the pore resembles that of the bulk solution, and at higher potentials, where the glycine concentration near the wall exceeds that of the bulk. In both scenarios, the impact is more pronounced for glycine dipeptides compared to peptides, as depicted in Figures 7 and 8. This enhanced effect can be attributed to a substantial increase in the local permittivity of the solution, which reduces walls’ charge screening by the mobile ions and intensifies electrostatic repulsion of the charged walls. In the case of relatively large surface potentials, where an increase in dipeptide glycine concentration leads to a significant increase in differential capacitance, as well as weakening of electrostatic screening in pores ( nm), the disjoining pressure increases due to the fact that the local glycine concentration within these pores is higher than in the surrounding solution. This leads to a higher osmotic pressure at the center of the pore compared to that in the bulk solution, not only due to the ionic contribution, as in the case of small potentials, but also due to the presence of excess osmolyte molecules.
@@ IV Discussion
Now, we would like to discuss possible implications of the predicted effects for practical applications.
First, it is particularly noteworthy that incorporating zwitterionic osmolytes into an electrolyte solution can lead to a substantial enhancement in differential capacitance, even at low electrolyte concentrations. This characteristic presents a significant advantage in electrochemical applications, as it effectively reduces unwanted electrochemical activity of ions on the electrode surface, without loss of differential capacitance. By partially substituting these ions with zwitterionic osmolytes the overall electrochemical performance of the system is improved. Their neutral charge prevents participation in Faradaic reactions, thereby reducing parasitic electrochemical activity (e.g., electrode corrosion) that typically plagues traditional electrolytes. This reduction in ionic activity could be beneficial as it minimizes adverse side reactions that can negatively affect the longevity and stability of the electrodes.
Second, during the operation of the supercapacitors, the micropores found within the electrodes are susceptible to collapse due to van der Waals forces acting between the pore walls, as well as chemical interactions with the ions present in the solution. Such structural degradation can severely compromise the performance and efficiency of the supercapacitor. In contrast, the osmolytic effects brought about by the zwitterionic osmolyte can enhance the repulsive forces acting between the pore walls, thereby reducing the likelihood of collapse and preserving the structural integrity of the micropores. Maintaining this integrity may be crucial, as it ensures that the porous electrodes retain a high surface area available for charge storage. A higher surface area correlates directly to improved capacitance and energy storage capabilities. Consequently, the introduction of zwitterionic osmolytes not only stabilizes the pore structure but also contributes to maintaining consistent capacitance levels over extended periods, ultimately enhancing the reliability and performance of supercapacitors in various applications. Thus, the strategic use of zwitterionic osmolytes can represent a significant advancement in electrolyte formulations, leading to more efficient energy storage devices.
@@ V Conclusion
In conclusion, we utilized modified Poisson-Boltzmann theory to systematically explore the role of zwitterionic osmolyte additives in modulating the disjoining pressure and differential capacitance of electric double layers within charged slit-like nanopores with conductive walls. Our findings revealed that increasing zwitterion concentration induces a pronounced synergistic enhancement of both disjoining pressure and differential capacitance, a dual effect critical for optimizing supercapacitor performance. We believe that the obtained insights may indicate on the unique potential of zwitterionic osmolytes as multifunctional additives for tailoring electric double layer properties, bridging the gap between capacitive efficiency and electrode durability. By elucidating the molecular mechanisms underlying this synergy, our work provides a foundational framework for designing advanced supercapacitors that harmonize high energy density with long-term operational stability. Future research could expand on these insights by exploring zwitterion-specific interactions in diverse electrolyte systems and nanopore geometries, further discovering their transformative potential for modern energy storage technologies.
@@ Acknowledgments
The authors thank the Russian Science Foundation (Grant No. 24-11-00096) for financial support.

### v3

@@ III.2 Disjoining pressure
Recent experimental studies by Govrin et al. [60] revealed that addition of the osmolytes to an electrolyte solution diminishes the electrostatic screening of charged surfaces by the mobile ions present in the solution, thereby amplifying the repulsive forces between them. In the context of slit-like pores with charged walls, this phenomenon is expected to result in an enhancement of the electrostatic component of the disjoining pressure as the concentration of zwitterionic molecules in the bulk solution increases. The MPB theory predicts this behavior, as illustrated in Figures 12(a,b). It suggests the rise of the disjoining pressure with increase of glycine concentration, both in its peptide and dipeptide form. Notably, this effect is observed at two distinct ranges of surface potentials: at low surface potentials ( in Fig.12a), where the glycine concentration within the pore resembles that of the bulk solution, and at higher potentials ( in Fig.12b), where the glycine concentration near the wall exceeds that of the bulk. In both scenarios, the impact is more pronounced for glycine dipeptides compared to peptides. This enhanced effect can be attributed to a substantial increase in the local permittivity of the solution, which reduces walls’ charge screening by the mobile ions and intensifies electrostatic repulsion of the charged walls. In the case of relatively large surface potentials, where an increase in dipeptide glycine concentration leads to a significant increase in differential capacitance, as well as weakening of electrostatic screening in pores, the disjoining pressure increases due to the fact that the local glycine concentration within these pores is higher than in the surrounding solution. This leads to a higher osmotic pressure at the center of the pore compared to that in the bulk solution, not only due to the ionic contribution, as in the case of small potentials, but also due to the presence of excess osmolyte molecules.
Finally, it would be reasonable to make a qualitative comparison with the mentioned experiment [60]. Strictly speaking, within present consideration we cannot directly compare our results to the experiment where the charge on silica surfaces was regulated due to cation desorption caused by zwitterion additives. To directly compare, it would be necessary to formulate a charge regulation model [49], which is outside the scope of this paper. Instead, we consider a hypothetical scenario of a spherical metal cantilever from an atomic force microscope interacting with a flat substrate of identical metal under the same surface potential in an aqueous electrolyte solution containing glycine in peptide form (see Fig.13). We assume that the electrostatic potential is sufficiently small to approximately correspond to the values of the surface charge density on the substrate accessible in Govrin et al.’s experiment. The force can be estimated using the Derjaguin’s approximation [13]
where represents the cantilever radius and represents the disjoining pressure embedded in the slit-like pore with a thickness of . As can be seen in Fig.14, an increase in the bulk zwitterion concentration from 0 to M results in a significant increase in the electrostatic repulsive force between the cantilever and the substrate, which is similar to what was observed in the experiments by Govrin et al. [60]
@@ IV Discussion
Now, we would like to speculate about possible implications of the predicted effects for practical applications. First, it is particularly noteworthy that incorporating zwitterionic osmolytes into an electrolyte solution can lead to a substantial enhancement in differential capacitance, even at low electrolyte concentrations. This characteristic presents a significant advantage in electrochemical applications, as it effectively reduces unwanted electrochemical activity of ions at the electrode surface, without loss of differential capacitance. By partially substituting these ions with zwitterionic osmolytes the overall electrochemical performance of the system is improved. Their neutral charge prevents participation in Faradaic reactions, thereby reducing parasitic electrochemical activity that typically plagues traditional electrolytes. This suppression of ionic activity may enhance electrode stability by mitigating parasitic side reactions that degrade its long-term performance. Second, during the operation of the supercapacitors, the micropores found within the electrodes are susceptible to collapse due to the electrode ageing [78, 79]. Such structural degradation can severely compromise the performance and efficiency of the supercapacitor. In contrast, zwitterionic osmolytes induce osmolytic effects that strengthen "intra-pore" repulsive forces, thereby reducing the likelihood of collapse and preserving the structural integrity of the micropores. Maintaining this integrity may be important, as it ensures that the porous electrodes retain a high surface area available for charge storage. A higher surface area correlates directly to improved capacitance and energy storage capabilities. Consequently, the introduction of zwitterionic osmolytes may not only stabilize the pore structure but also contribute to maintaining consistent capacitance levels over extended periods, ultimately enhancing the reliability and performance of supercapacitors in various applications.
Finally, it is instructive to discuss the possible limitations of present theoretical model. It is important to note that this model neglects the electrostatic correlations between ions and polar molecules of glycine and water. When implicitly accounting for the solvent using a constant permittivity of the solvent medium, it is known that correlation effects, including those arising from image forces, can be addressed within the framework of variational field theory through the electrostatic Green’s function [33, 52, 32, 53]. However, when explicitly considering the polar solvent or co-solvent molecules, solving the equations of variational field theory becomes significantly more challenging. Tackling this complex numerical problem will allow for a more precise refinement of the effects predicted in this study, which is certainly a topic for further research.
@@ V Conclusion
In conclusion, we utilized modified Poisson-Boltzmann theory to systematically explore the role of zwitterionic osmolyte additives in modulating the disjoining pressure and differential capacitance of electric double layers within charged slit-like nanopores with conductive walls. Our findings revealed that increase of zwitterion concentration induces a pronounced synergistic enhancement of both disjoining pressure and differential capacitance, a dual effect potentially useful for optimizing supercapacitor performance. We believe that our findings reveal the unique potential of zwitterionic osmolytes as multifunctional additives capable of tailoring electric double layer properties, bridging the gap between capacitive efficiency and electrode durability. By elucidating the molecular mechanisms underlying this synergy, our work provides a foundational framework for designing advanced supercapacitors that harmonize high energy density with long-term operational stability.
@@ Acknowledgments
The authors thank the Russian Science Foundation (Grant No. 24-11-00096) for financial support of parts of this paper devoted to the theoretical background and disjoining pressure calculations. The part on the osmolyte’s effect on the differential capacitance was partly supported by the Ministry of Science and Higher Education of the Russian Federation.